iamdead:

A nut mixture of almonds and macadamia nuts at a small fair is $1.00 per pound of almonds and $5.93 per pound of macadamia nuts. Over the entire day, 102 pounds of the nut mixture were sold for $452.03. If p is the number almonds and n is the number of macadamia nuts, then the system of equations that models this scenario is: p+n=102p+5.93n=452.03 Determine the correct description and amount of pounds for almonds and macadamia nuts that were sold. A (61.0, 41.0) There were 4.0 pounds of almonds and 36.0 pounds of macadamia nuts sold at the fair. B (31.0, 71.0) There were 31.0 pounds of almonds and 71.0 pounds of macadamia nuts sold at the fair. C (41.0, 61.0) There were 36.0 pounds of almonds and 4.0 pounds of macadamia nuts sold at the fair. D (71.0, 31.0) There were 71.0 pounds of almonds and 31.0 pounds of macadamia nuts sold at the fair.

Vocaloid:

p+n=102 p+5.93n=452.03 is your system you can subtract the two equations (2nd eqn minus 1st eqn) to get p - p = 0 5.93n - n = 4.93n 452.03 - 102 = 350.03 meaning 4.93n = 350.03 solve for n (the pounds of macadamia nuts) and choose the right solution based on this gtg